OpenStudy (anonymous):

solve by factoring 4x^2 -5x +1=0

OpenStudy (anonymous):

Do you know how to factor?

OpenStudy (anonymous):

4x^2 -5x +1=0 4x^2 -4x -x +1=0 4x(x-1) -1(x-1) =0 (4x-1)(x-1)=0 and so on..